Problem:
Two angles are supplementary. Three times the smaller is five less than twice the larger. Find the number of degrees in each angle.
: Problem:
Two angles are supplementary. Three times the smaller is five less than twice the larger. Find the number of degrees in each angle.

Answer by ankor@dixie-net.com(4484)

Two angles are supplementary.
:
let x = 1 angle
then
(180-x) = it's supplement
:
"Three times the smaller is five less than twice the larger."
3x = 2(180-x) - 5
3x = 360 - 2x - 5
3x + 2x = 360 - 5
5x = 355
x = 355/5
x = 71
:
Find the number of degrees in each angle.
71 degrees = smaller angle
:
180 - 71 = 109 degrees is the larger angle
:
:
Check solutions in the statement:
"Three times the smaller is five less than twice the larger."
3(71) = 2(109) - 5
213 = 218 - 5